# Contrastive Self-Supervised Learning As Neural Manifold Packing

### v1

@@ Abstract
Contrastive self-supervised learning based on point-wise comparisons has been widely studied for vision tasks. In the visual cortex of the brain, neuronal responses to distinct stimulus classes are organized into geometric structures known as neural manifolds. Accurate classification of stimuli can be achieved by effectively separating these manifolds, akin to solving a packing problem. We introduce Contrastive Learning As Manifold Packing (CLAMP), a self-supervised framework that recasts representation learning as a manifold packing problem. CLAMP introduces a loss function inspired by the potential energy of short-range repulsive particle systems, such as those encountered in the physics of simple liquids and jammed packings. In this framework, each class consists of sub-manifolds embedding multiple augmented views of a single image. The sizes and positions of the sub-manifolds are dynamically optimized by following the gradient of a packing loss. This approach yields interpretable dynamics in the embedding space that parallel jamming physics, and introduces geometrically meaningful hyperparameters within the loss function. Under the standard linear evaluation protocol, which freezes the backbone and trains only a linear classifier, CLAMP achieves competitive performance with state-of-the-art self-supervised models. Furthermore, our analysis reveals that neural manifolds corresponding to different categories emerge naturally and are effectively separated in the learned representation space, highlighting the potential of CLAMP to bridge insights from physics, neural science, and machine learning.
@@ 1 Introduction
Learning image representations that are both robust and broadly transferable remains a major challenge in computer vision, with critical implications for the efficiency and accuracy of training across downstream tasks. State-of-the-art models address this problem by employing contrastive self-supervised learning (SSL), wherein embeddings are treated as vectors and optimized using pairwise losses to draw together positive samples, augmented views of the same image, and push apart negative samples,views of different images [1, 2, 3, 4, 5, 6]. While SSL has advanced significantly, and in some cases, surpassed supervised methods [7, 8], its geometric foundations remain largely underexplored. In visual cortex, neural activity often resides within geometric structures known as neural manifolds — which span the embedding space [9] and may exhibit low intrinsic dimensionality [10] that capture much of the observed variability in neural activity. Neural manifolds have been characterized in a number of neural systems supporting a variety of cognitive processes (in addition to vision) including neural activity coordination, motor control, and information coding [11, 12, 13, 14]. In the context of SSL, neural representations of similar and dissimilar samples are frequently conceptualized as distinct sub-manifolds in the representation space, where a sub-manifold corresponds to the encoding of a single data sample, and a manifold to an entire class. In classification, the objective is to separate these class-level manifolds so that each maps cleanly to a distinct category under a readout function. When manifolds corresponding to different classes, for example, manifolds representing “cats” and “dogs” overlap in neural state space, the resulting ambiguity can lead to misclassification. This viewpoint recasts classification as the problem of efficiently arranging manifolds in representation space to minimize overlap and enhance separability [15].
@@ 2.1 Interacting particle models in physics
A particularly insightful analogy for CLAMP comes from random organization models, originally introduced to describe the dynamics of driven colloidal suspensions [16, 17, 18]. In these models, overlapping particles are randomly “kicked” while isolated particles remain static. As the volume fraction is increased, the system undergoes a phase transition from an absorbing state (where no particles overlap) to an active steady state (where overlaps cannot be resolved under the prescribed dynamics) in two and three dimensions [19]. When kicks between overlapping particle pairs are reciprocal, meaning that each particle is displaced by an equal but opposite amount, the critical and active states are characterized by anomalously suppressed density fluctuations at large scales, a property known as “hyperuniformity”[20] or “blue noise”[21, 22]. This absorbing-to-active transition coincides with random close packing (RCP) when the noisy dynamics are reciprocal [17]. Recent work further establishes an equivalence between this process and a stochastic update scheme that minimizes a repulsive energy under multiplicative noise [15]. We note that even beyond the hyperuniformity observed for dense active states of random organization systems, for a high-dimensional system of short-range, repulsive particles, such as those interacting via an exponential potential ($e^{-cr},c>0$), confined to a closed surface, a uniform spatial distribution minimizes the system’s energy in the large density limit [23, 24, 25]. Drawing on these insights, CLAMP employs a short-range repulsive potential over augmentation sub-manifolds as its self-supervised learning loss, enhancing embedding separation with only local operations.
@@ 2.2 Feature-level contrastive learning
Following recent SOTA self-supervised learning models, we use image augmentations and contrastive losses. Most of these self-supervised learning frameworks treat features as points or vectors in the embedding space, contrasting positive samples—embeddings of augmented views of the same image—with negative samples—embeddings of augmented views of different images—using a pairwise contrastive loss such as InfoNCE [1, 2, 3, 4, 5, 6]. Similarly to previous approaches, CLAMP treats each image as its own class; however, instead of applying the contrastive loss directly to individual feature vectors, it operates on the entire augmentation sub-manifolds surrounding each input embedding.
@@ 2.3 Clustering methods for self-supervised learning
Clustering-based self-supervised learning methods assign images to clusters based on similarity criteria, rather than treating each image as a distinct class. For instance, DeepCluster [26] applies the k-means algorithm to assign pseudo-labels to embeddings, which serve as training targets in subsequent epochs. However, this approach requires k-means classification on the entire dataset, making it unsuitable for online, mini-batch, learning. SwAV [27] overcomes this limitation by using soft assignments and jointly learning cluster prototypes and embeddings, enabling flexible online updates. In contrast to clustering methods, our approach groups each image with its augmentations to learn the geometry of the augmentation sub-manifold without clustering across different samples.
@@ 2.4 Alignment-Uniformity and maximum manifold capacity methods
The Alignment-Uniformity method [28] constructs the loss function by incorporating two components: an alignment term and a uniformity term. The alignment term maximizes the cosine similarity of the embedding vectors between positive pairs, thereby encouraging local consistency. The uniformity term acts as a repulsive pairwise potential over all pairs of embeddings, driving the representations to be uniformly distributed on the hyper-sphere and mitigating representational collapse [29].
@@
Unlike MMCR, which expressly maximizes the representation’s binary linear classification capacity, CLAMP optimizes the embeddings for an $n$-ary nonlinear classification problem, making the two approaches fundamentally different. Moreover, CLAMP’s loss function is based on local interactions: the contribution from each augmentation sub-manifold depends only on its neighboring sub-manifolds, whereas MMCR requires global information from all sub-manifolds. Finally, although MMCR has comparable computational complexity to CLAMP, MMCR is less efficient to train, because it requires one SVD for each learning step, compared to CLAMP, and its loss function and optimization dynamics lack the clear, physics-based interpretability that CLAMP provides.
@@ 3.1 Description
We employ a backbone encoder $f_{\theta}(x)$ as the feature extractor to get the representations, $y_{i}^{k}=f_{\theta}(x_{i}^{k})$. To map these representations into the embedding space, we apply a multilayer perceptron (MLP) projection head, $z_{i}^{k}=g_{\theta}(y_{i}^{k})$. Normalizing embedding vectors onto the unit hypersphere has been shown to be effective in self-supervised learning and variational encoders [32, 28, 34]. Therefore, we normalize the embeddings to lie on the unit hypersphere by subtracting the global center and projecting to unit norm, $\tilde{z}_{i}^{k}=(z_{i}^{k}-c)/||z_{i}^{k}-c||_{2}$, where $c=\frac{1}{mb}\sum_{i}\sum_{k}z_{i}^{k}$ is the global mean of all the projected vectors in the batch.
@@
$\begin{split}&\mathcal{L}_{overlap}=\begin{cases}\sum\limits_{i\neq j}\left(1-\frac{\lVert Z_{i}-Z_{j}\rVert_{2}}{r_{i}+r_{j}}\right)^{2}\text{,if }\lVert Z_{i}-Z_{j}\rVert_{2}<r_{i}+r_{j}\\ 0\text{,\quad otherwise }\end{cases},\text{where }r_{i}=r_{s}\sqrt{\frac{\mathrm{Tr}(\Lambda_{i})}{b}}.\end{split}$ (1)
Since the logarithm is a monotonic function, applying it to $\mathcal{L}_{\text{overlap}}$ preserves the locations of its minima while compressing large loss values, stabilizing the training process.
Following the convention established in [2], we refer the output space of the encoder $f_{\theta}(\cdot)$ as the representation space, while the output space of the projection head $g_{\phi}(f_{\theta}(\cdot))$ is designated as the embedding space.
@@ 3.1.1 Approximate augmentation sub-manifolds as ellipsoids
We approximate the distribution of embedding vectors $p(\tilde{z}_{i})$ generated from random augmentations of input image $x_{i}$ as a multivariate Gaussian with mean $\tilde{Z}_{i}$ and covariance $\Lambda_{i}$, $\tilde{z}_{i}\sim\mathcal{N}(Z_{i},\Lambda_{i})$. We assume that $\Lambda_{i}$ is full-rank and invertible (see Appendix J for a treatment of the general case). To represent the augmentation sub-manifold in embedding space, we define the corresponding enclosing ellipsoid with Mahalanobis distance $r_{s}$,
@@
where $r_{s}$ is a scaling hyperparameter that controls the effective size of the manifold. A large $r_{s}$ corresponds to larger ellipsoids, and an appropriate choice of $r_{s}$ ensures that the ellipsoid encloses the majority of the augmentation sub-manifold (Fig. 2(a)). The lengths of the ellipsoid’s semi-axes are given by the square roots of the nonzero eigenvalues of $\Lambda_{i}$ scaled by $r_{s}$: $r_{s}\sqrt{\lambda_{i}^{(1)}},r_{s}\sqrt{\lambda_{i}^{(2)}},...$ where $\lambda_{i}^{(k)}$ is the k-th eigenvalue of $\Lambda_{i}$. To estimate the effective size of the ellipsoid, we compute the square root of the expected squared semi-axial length, $\sqrt{\mathbb{E}[(r_{s}\sqrt{\lambda})^{2}]}$, yielding the sub-manifold radius, $r_{i}=r_{s}\sqrt{\frac{\sum_{j}\lambda_{i}^{(j)}}{\text{rank}(\Lambda_{i})}}$. The radius $r_{i}$ provides an upper bound on the volume of the ellipsoid, $r_{i}\geq cV_{i}^{\min(b,D)}$ (see Appendix J for details). In practice, the number of augmentations $m\sim O(10)$ is much smaller than the embedding dimension $D\sim O(100)$, making the empirical covariance $\Lambda_{i}$ lower rank or singular. In such cases, the expression for $r_{i}$ remains valid with $\mathrm{rank}(\Lambda_{i})\approx m$ serving as a proxy (see Appendix J for discussion).
@@
CLAMP models sub-manifolds as repulsive particles interacting via a short-range potential $E(Z_{i},Z_{j})$, where $Z_{i},Z_{j}$ are the centroids of sub-manifolds $i$ and $j$. The potential attains its maximum at $||Z_{i}-Z_{j}||_{2}=0$ corresponding to complete overlap between two sub-manifolds and drops to zero when $||Z_{i}-Z_{j}||_{2}>r_{i}+r_{j}$, ensuring that non-overlapping sub-manifolds do not interact. Eq. 1 implies that increasing the inter-center distance between sub-manifolds and reducing their sizes leads to a lower loss.
@@
Separability. Separability encourages the negative embeddings to be distant and separable on the unit hypersphere. Analogous repulsive interactions in physical systems are known to produce near-uniform distributions. Consistent with this, we observe that the embeddings progressively become more separable throughout training. To quantify this effect we track both similarity (distance among positives) and separability (distance among negatives) throughout pretraining, confirming that CLAMP systematically enhances class-level separation.
@@ 3.2.3 Optimization
We used the LARS optimizer [39] for training the network. For CIFAR10, we trained the model for 1000 epochs using the warmup-decay learning rate scheduler, and with 10 warmup steps and a cosine decay for rest of the steps. For ImageNet-100, we trained the model for 200 epochs with 10 steps of warmup and 190 steps of cosine decay for the learning rate. For ImageNet-1K, we trained the model for 100 epochs with 10 steps of warmup and 90 steps of cosine decay for the learning rate. Training on 8 A100 GPUs with batch size 1024 for 4 views with distributed data parallelization for 100 epochs took approximately 17 hours. See Appendix D for details about the hyperparameter and the training setup.
Linear evaluation Semi-supervised learning
@@
MMCR (4 views)[32] 82.88 71.5 49.4 66.0
  CLAMP (4 views) 83.96 ±.$\pm.19$ 69.16 ±.$\pm.18$ 46.7 ±.$\pm.15$ 64.6 ±.$\pm.12$
Table 1: Linear evaluation accuracy on ImageNet-100 dataset for 200-epoch pretraining: For SimCLR the higher accuracy between [32] and [41] is reported. For Barlow twins, BYOL, and MoCo-V2, the accuracies using ResNet-18 as the backbone (labeled with *) in [41] are reported because either ResNet-50 results are absent or ResNet-18 performs better in [41]. A&U and MMCR results are taken from [28] and [32] respectively. Linear evaluation accuracy on ImageNet-1K dataset for 100-epoch pretraining: For SimCLR, SWAV, Barlow Twins, BYOL, VICReg, CorInfoMax, MoCo-V2 and SimSiam, we reported the highest accuracy for each method among the results in the solo-learn library [42], VISSL library[43], [32] and [41]. We report the 200-epoch pretraining result for A&U, as the 100-epoch result is unavailable [28]. Semi-superivsed learning:The semi-supervised learning results for VICReg and CorInfoMax are taken from [41], those for other methods are from [32].

### v2

@@ Abstract
Contrastive self-supervised learning based on point-wise comparisons has been widely studied for vision tasks. In the visual cortex of the brain, neuronal responses to distinct stimulus classes are organized into geometric structures known as neural manifolds. Accurate classification of stimuli can be achieved by effectively separating these manifolds, akin to solving a packing problem. We introduce Contrastive Learning As Manifold Packing (CLAMP), a self-supervised framework that recasts representation learning as a manifold packing problem. CLAMP introduces a loss function inspired by the potential energy of short-range repulsive particle systems, such as those encountered in the physics of simple liquids and jammed packings. In this framework, each class consists of sub‑manifolds embedding multiple augmented views of a single image. The sizes and positions of the sub-manifolds are dynamically optimized by following the gradient of a packing loss. This approach yields interpretable dynamics in the embedding space that parallel jamming physics, and introduces geometrically meaningful hyperparameters within the loss function. Under the standard linear evaluation protocol, which freezes the backbone and trains only a linear classifier, CLAMP achieves competitive performance with state‑of‑the‑art self‑supervised models. Furthermore, our analysis reveals that neural manifolds corresponding to different categories emerge naturally and are effectively separated in the learned representation space, highlighting the potential of CLAMP to bridge insights from physics, neural science, and machine learning.
@@ 1 Introduction
Learning image representations that are both robust and broadly transferable remains a major challenge in computer vision, with critical implications for the efficiency and accuracy of training across downstream tasks. State-of-the-art models address this problem by employing contrastive self-supervised learning (SSL), wherein embeddings are treated as vectors and optimized using pairwise losses to draw together positive samples, augmented views of the same image, and push apart negative samples, views of different images [1, 2, 3, 4, 5, 6]. While SSL has advanced significantly, and in some cases, surpassed supervised methods [7, 8], its geometric foundations remain largely underexplored. In visual cortex, neural activity often resides within geometric structures known as neural manifolds — which span the embedding space [9] and may exhibit low intrinsic dimensionality [10] that capture much of the observed variability in neural activity. Neural manifolds have been characterized in a number of neural systems supporting a variety of cognitive processes (in addition to vision) including neural activity coordination, motor control, and information coding [11, 12, 13, 14]. In the context of SSL, neural representations of similar and dissimilar samples are frequently conceptualized as distinct sub-manifolds in the representation space, where a sub-manifold corresponds to the encoding of a single data sample and its transformations, and a manifold to an entire class. In classification, the objective is to separate these class-level manifolds so that each maps cleanly to a distinct category under a readout function. When manifolds corresponding to different classes, for example, manifolds representing “cats” and “dogs” overlap in neural state space, the resulting ambiguity can lead to misclassification. This viewpoint recasts classification as the problem of efficiently arranging manifolds in representation space to minimize overlap and enhance separability [15].
@@ 2.1 Interacting particle models in physics
A particularly insightful analogy for CLAMP comes from random organization models, originally introduced to describe the dynamics of driven colloidal suspensions [16, 17, 18]. In these models, overlapping particles are randomly “kicked” while isolated particles remain static. As the volume fraction is increased, the system undergoes a phase transition from an absorbing state (where no particles overlap) to an active steady state (where overlaps cannot be resolved under the prescribed dynamics) in two and three dimensions [19]. When kicks between overlapping particle pairs are reciprocal, meaning that each particle is displaced by an equal but opposite amount, the critical and active states are characterized by anomalously suppressed density fluctuations at large scales, a property known as “hyperuniformity”[20] or “blue noise”[21, 22]. This absorbing-to-active transition coincides with random close packing (RCP) when the noisy dynamics are reciprocal [17]. Recent work further establishes an equivalence between this process and a stochastic update scheme that minimizes a repulsive energy under multiplicative noise [15]. We note that even beyond the hyperuniformity observed for dense active states of random organization systems, for a high-dimensional system of short-range, repulsive particles, such as those interacting via an exponential potential ($e^{-cr},c>0$), confined to a closed surface, a uniform spatial distribution minimizes the system’s energy in the large density limit [23, 24, 25]. Drawing on these insights, CLAMP employs a short-range repulsive potential over augmentation sub-manifolds as its self-supervised learning loss, favoring the separation of embeddings.
@@ 2.2 Feature-level contrastive learning
Following recent SOTA self-supervised learning models, we use image augmentations and contrastive losses. Most self-supervised learning frameworks treat features as points or vectors in the embedding space, contrasting positive samples—embeddings of augmented views of the same image—with negative samples. They typically use a pairwise contrastive loss such as InfoNCE [1, 2, 3, 4, 5, 6], or a weighted-sample InfoNCE loss where negative samples are drawn from tailored Boltzmann-like distributions to sharpen sample-wise contrasts [26]. Here, CLAMP treats each image as its own class; however, instead of applying the contrastive loss directly to individual feature vectors, it operates on the entire augmentation sub-manifolds surrounding each input embedding.
@@ 2.3 Clustering methods for self-supervised learning
Clustering-based self-supervised learning methods assign images to clusters based on similarity criteria, rather than treating each image as a distinct class. For instance, The DeepCluster [27] applies the k-means algorithm to assign pseudo-labels to embeddings, which serve as training targets in subsequent epochs. Collapse-resistant non-contrastive methods prevent feature collapse by using a fixed dictionary and enforcing augmentation invariance without negative samples [28]. However, the DeepCluster approach requires k-means classification on the entire dataset, making it unsuitable for online, mini-batch, learning. SwAV [29] overcomes this limitation by using soft assignments and jointly learning cluster prototypes and embeddings, enabling flexible online updates. In contrast to clustering methods, our approach groups each image with its augmentations to learn the geometry of the augmentation sub-manifold without clustering across different samples.
@@ 2.4 Alignment-Uniformity and maximum manifold capacity methods
The Alignment-Uniformity method [30] constructs the loss function by incorporating two components: an alignment term and a uniformity term. The alignment term maximizes the cosine similarity of the embedding vectors between positive pairs, thereby encouraging consistency. The uniformity term acts as a repulsive pairwise potential over all pairs of embeddings, driving the representations to be uniformly distributed on the hyper-sphere and mitigating representational collapse [31].
@@
Unlike MMCR, which expressly maximizes the representation’s binary linear classification capacity, CLAMP optimizes the embeddings for an $n$-ary nonlinear classification problem, making the two approaches fundamentally different. Moreover, CLAMP’s loss function is based on short-range interactions: the contribution from each augmentation sub-manifold depends only on its neighboring sub-manifolds, whereas MMCR requires global information from all sub-manifolds. Finally, although MMCR has comparable computational complexity to CLAMP, MMCR is less efficient to train, because it requires one SVD for each learning step, compared to CLAMP, and its loss function and optimization dynamics lack the clear, physics-based interpretability that CLAMP provides.
@@ 3.1 Description
[…]
@@
$\begin{split}&\mathcal{L}_{overlap}=\begin{cases}\sum\limits_{i\neq j}\left(1-\frac{\lVert Z_{i}-Z_{j}\rVert_{2}}{r_{i}+r_{j}}\right)^{2}\text{, if }\lVert Z_{i}-Z_{j}\rVert_{2}<r_{i}+r_{j}\\ 0\text{,\qquad otherwise }\end{cases},\text{where }r_{i}=r_{s}\sqrt{\frac{\mathrm{Tr}(\Lambda_{i})}{m}}.\end{split}$ (1)
Note that even though the loss function acts locally on neighboring sub-manifolds in the embedding space, each weight update affects all embeddings. Since the logarithm is a monotonic function, applying it to $\mathcal{L}_{\text{overlap}}$ preserves the locations of its minima while compressing large loss values, stabilizing the training process.
Following the convention established in [2], we refer to the output space of the encoder $f_{\theta}(\cdot)$ as the representation space, while the output space of the projection head $g_{\phi}(f_{\theta}(\cdot))$ is designated as the embedding space.
@@ 3.1.1 Approximate augmentation sub-manifolds as ellipsoids
We approximate the distribution of embedding vectors $p(\tilde{z}_{i})$ generated from random augmentations of input image $x_{i}$ as a multivariate Gaussian with mean $\tilde{Z}_{i}$ and covariance $\Lambda_{i}$, $\tilde{z}_{i}\sim\mathcal{N}(Z_{i},\Lambda_{i})$. We assume that $\Lambda_{i}$ is full-rank and invertible (see Appendix K for a treatment of the general case). To represent the augmentation sub-manifold in embedding space, we define the corresponding enclosing ellipsoid with Mahalanobis distance $r_{s}$,
@@
where $r_{s}$ is a scaling hyperparameter that controls the effective size of the manifold. A large $r_{s}$ corresponds to larger ellipsoids, and an appropriate choice of $r_{s}$ ensures that the ellipsoid encloses the majority of the augmentation sub-manifold (Fig. 2(a)). The lengths of the ellipsoid’s semi-axes are given by the square roots of the nonzero eigenvalues of $\Lambda_{i}$ scaled by $r_{s}$: $r_{s}\sqrt{\lambda_{i}^{(1)}},r_{s}\sqrt{\lambda_{i}^{(2)}},$... where $\lambda_{i}^{(k)}$ is the k-th eigenvalue of $\Lambda_{i}$. To estimate the effective size of the ellipsoid, we compute the square root of the expected squared semi-axial length, $\sqrt{\mathbb{E}[(r_{s}\sqrt{\lambda})^{2}]}$, yielding the sub-manifold radius, $r_{i}=r_{s}\sqrt{\frac{\sum_{j}\lambda_{i}^{(j)}}{\text{rank}(\Lambda_{i})}}$. The radius $r_{i}$ provides an upper bound on the volume of the ellipsoid $i$ ($V_{i}$), $r_{i}\geq cV_{i}^{1/\min(b,D)}$ (see Appendix K for details). In practice, the number of augmentations $m\sim O(10)$ is much smaller than the embedding dimension $D\sim O(100)$, making the empirical covariance $\Lambda_{i}$ lower rank or singular. In such cases, the expression for $r_{i}$ remains valid with $\mathrm{rank}(\Lambda_{i})\approx m$ serving as a proxy (see Appendix K for discussion).
@@
CLAMP models sub-manifolds as repulsive particles interacting via a short-range potential $E(Z_{i},Z_{j})$, where $Z_{i},Z_{j}$ are the centroids of sub-manifolds $i$ and $j$. The potential attains its maximum at $Z_{i}-Z_{j}||_{2}=0$ corresponding to complete overlap between two sub-manifolds and drops to zero when $Z_{i}-Z_{j}||_{2}>r_{i}+r_{j}$, ensuring that non-overlapping sub-manifolds do not interact. Eq. 1 implies that increasing the inter-center distance between sub-manifolds and reducing their sizes leads to a lower loss.
@@
Separability. Separability encourages the negative embeddings to be distant and separable on the unit hypersphere. Analogous repulsive interactions in physical systems are known to produce near-uniform distributions. Consistent with this, we observe that the embeddings progressively become more separable throughout training. To quantify this effect, we track both similarity (distance among positives) and separability (distance among negatives) throughout pretraining, confirming that CLAMP systematically enhances class-level separation.
@@ 3.2 Implementation details
The code for CLAMP is available at https://github.com/guanming-zhang/clamp
@@ 3.2.3 Optimization
We used the LARS optimizer [41] for training the network. For CIFAR10, we trained the model for 1000 epochs using the warmup-decay learning rate scheduler, and with 10 warmup steps and a cosine decay for rest of the steps. For ImageNet-100, we trained the model for 200 epochs with 10 steps of warmup and 190 steps of cosine decay for the learning rate. For ImageNet-1K, we trained the model for 100 epochs with 10 steps of warmup and 90 steps of cosine decay for the learning rate. Training on 8 A100 GPUs with batch size 1024 for 4 views with distributed data parallelization for 100 epochs took approximately 17 hours. See Appendix D for details.
Linear evaluation Semi-supervised
@@
MMCR (4 views+ME)[34] 82.88 71.5 49.4 66.0
  CLAMP (4 views) 85.12 $\pm.05$ 69.50 $\pm.14$ 47.38 $\pm.56$ 65.10 $\pm.30$
CLAMP (8 views) 85.10 $\pm.15$ 70.04 $\pm.16$ 47.87 $\pm.03$ 65.96 $\pm.04$
Table 1: Linear evaluation accuracy on ImageNet-100 dataset for 200-epoch pretraining: For SimCLR the higher accuracy between [34] and [43] is reported. For Barlow twins, BYOL, and MoCo-V2, the accuracies using ResNet-18 as the backbone (labeled with *) in [43] are reported because either ResNet-50 results are absent or ResNet-18 performs better in [43]. A&U and MMCR results are taken from [30] and [34] respectively. Linear evaluation accuracy on ImageNet-1K dataset for 100-epoch pretraining: For SimCLR, SWAV, Barlow Twins, BYOL, VICReg, CorInfoMax, MoCo-V2 and SimSiam, we reported the highest accuracy for each method among the results in the solo-learn library [44], VISSL library [45], [34] and [43]. We report the 200-epoch pretraining result for A&U, as the 100-epoch result is unavailable [30]. Semi-supervised learning:The semi-supervised learning results for VICReg and CorInfoMax are taken from [43], those for other methods are from [34]. In MMCR with 4 views and a momentum encoder (4 views + ME), the effective 8 views, 4 from the backbone and 4 from the encoder, achieve the same linear accuracy (71.5%) as using 8 views without the momentum encoder [34]